point which of the following is equal to the integral w+) + 1) Infxº)dx + A(+1)- 2*(x2 + 1)(x In x - x) dx. •(3+x) (3 inc?)) is this correct ?
Answer #1

+ (x) 3(x) do Sfga fg - felg f = ln (x²) ; g=x+ fs 9.*** ad id (x²)=nan + 1 $x? dx = x dx = xht te n+) (2x) 2 ts a da -)(